Members of the Computer Club were assessed equal amounts to raise $1200 to buy some software. When 8 new members joined, the per-member assessment was reduced by $7.50. What was the new size of the club? Thank you.
Answer by Paul(988)

Current members = x
8 new memebers enetered the current = x+8
7.50(x+x+8)=1200
15x+60=1200
15x=1140
x=76
The current memebrs were 76.
76+8=84

Hence, when 8 people joined the memeber in club are 84 people.
Paul.